OpenStudy (anonymous):

if you use the power rule to find the derivative of (7x^2-2) Is it just 14x? what happens to the -2? or can u not use the power rule because of the -2?

OpenStudy (valpey):

Yes. The derivative of a constant is zero.

OpenStudy (campbell_st):

well if you think of \[f(x) = -2x^0\] then applying the derivative rule will yield \[f'(x) = 0 \times (-2x^{-1}) = 0\]

OpenStudy (campbell_st):

it can be written this way as \[-2\times x^0 = -2 \times 1 = -2\]

OpenStudy (anonymous):

Remember that when you add two functions, you can add their respective derivatives. \[\frac{d}{dx}(7x^2-2) = \frac{d}{dx}(7x^2) + \frac{d}{dx}(-2) = \frac{d}{dx}(7x^2) + 0 = \frac{d}{dx}(7x^2) \] Then we use the power rule: \[\frac{d}{dx}(7x^2) = 14x \]

OpenStudy (anonymous):

would you ever subtract the two functions, rather then add? or do you always add and apply a negative value to the second function?

OpenStudy (campbell_st):

it applies to any constant f(x) = 3x^2 + 11 you have f'(x) = 6x one of the 1st concepts to understand in differentiation for any constant f(x) = c then f'(x) = 0

OpenStudy (anonymous):

thanks for your help guys! :)